An element of R is called idempotent if a 2=a. Show that set of all idempotents in a commutative ring is closed under multiplication.


1
Expert's answer
2021-09-06T19:11:45-0400

Suppose that "a,b\in R" are both idempotent. Let us prove that "ab" is idempotent (which is equivalent to the fact that "ba" is idempotent by symmetry of choice of a and b). We will verify it directly :

"(ab) \cdot (ab)=abab", by commutativity we have "(ab)\cdot (ab)=a^2b^2", which by idempotency of a and b gives us "(ab)\cdot(ab)=ab", so it is idempotent. Therefore, the set of idempotent elements is closed under multiplication.